UT 08 Apr 2022 - DEC Night



Night Info

  • Sunset 18:44
  • Nautical Twilight Ends 19:39
  • Astro Twilight Ends 20:09
  • Moon sets 01:11
  • Astro Twilight Starts 04:34
  • Nautical Twilight Starts 05:04
  • Sunrise 05:59

Staff

  • OSA: D.G.Huerta
  • ISp: J.Schuh
  • Mtn Manager: J.Prothro
  • ISA/Observer: O.Kuhn, A Becker
  • Special Guests:

Night Plan

  • Starting Configuration LBCs
    • Please do the same as you did last night. Start active TMS at zenith before the dome is opened at leave it running while shutters opened. JMH thinks he fixed the primary limit problem.
    • PHAs w/TMS (no guiding) - 3hr (20min each)
      • 2017 VW13
      • 2018 LB1
      • 2019 XO2
      • Cycle through the three scripts, hopefully thrice (three times)
    • LBC Extra Focal images in equilibrium -IT 8608, IT 8592 - 1hr
      • telescope is in good equilibrium, collimate and then inject Z4/Z11 after which we would take extra-focal (dZ=-0.8) images with offsets to Z22 and Z11
      • seeing < 1.2", any clouds
  • Reconfigure for Bino-LUCI - 00:29-00:59
  • LUCI-1 AO/ESM on-sky checks - 01:00-02:00
    • OT files: /home/lbto/Service_Observing/2022A_D_E/LUCI/AO_ESM/
      • AO V84, ESM V37
    • OSCO - 0.5hr
    • Seeing < 1.5" for ESM - 0.5hr
      • Bin 2 ESM railed camera test
      • Railed on Mar 27 - UT 05:18 (local 22:18) - PA 207.1 Parallactic -164.67
      • OT: /home/lbto/Service_Observing/2022A_D_E/LUCI/AO_ESM/ESM_V38.xml
        • Folder: Bin 2 Railed Camera Test
      • Select the script closest in local time to the observations to match the rotator angle with the problem angle from before
  • Reconfigure for MODS-Bino - 02:00-02:15
  • GRB 180618A Followup - 1.25hr w/overheads - 02:15-03:30
    • MODS dual grating
    • Seeing <1.3", thin clouds - if clouds are thicker, take an extra 1200 sec exposure
    • OT: /home/lbto/Service_Observing/2022A_D_E/MODS/Rossi/grb180618A_mods-ot_V2.xml
    • Finder Chart: fc-slit.png
  • V1674 Her Followup - MODS Bino - M.Wagner - 03:30-04:15 0.75hr
    • Visible 02hr-twilight
    • seeing < 1.8" and clouds 1-2mag
    • OT Files and finder chart: /home/lbto/Service_Observing/2022A_D_E/MODS/MWagner/V1674Her
  • Spectrophotometric Standard - 04:15-04:30
    • BD +33 2642 Dual Grating

Poor Weather Backup

  • LBC Photomeric Standards - requires clear skies (during 1st half)
  • LUCI Photometric Standards (any seeing, if so bad can't guide, use track mode) - requires clear skies
    • /home/lbto/Service_Observing/2022A_D_E/PoorWeather/LUCI_Phot_Standards
    • Select stars based on visibility, need to get stars at two different airmasses
  • MODS Spectrophotometric Stds and Photometric Standards - requires clear skies
  • Poor Weather Science Backup - LUCI BINO
    • Rothberg - Local Mergers
    • OT: /home/lbto/Service_Observing/2022A_D_E/PoorWeather/Rothberg/LocalMergers/COBandhead_Mergers_V42.xml
    • Not in order of priority
      • NGC 4004 - 21hr-twilight Dec +27
      • NGC 3597 - 23hr-02hr Dec -23
      • NGC 6598 - 03hr-twilight Dec +69
      • UGC 9829 - 01:30hr-twilight Dec -01
      • UGC 10607 - 02hr -twilight Dec +26
      • UGC 10675 - 02hr-twilight Dec +31

Night Log

01:47 I copied the TMS reference from the previous night, 20220407, to /tmp and started the active loop. We are at zenith, waiting to open.

01:50 David is slewing to 30 deg to open the shutter doors. TMS has been sending updates at 30 deg, and the mirrors have not gone into the limits.

01:59 But there has to be an alarm... the right shutter door is not opening. David and Mike are working on it. They have called Bruce.

02:02 Like last night, the total coeffs for astigmatism on SX are Z5=-900 Z6=-110 nm. Z11=1000 as expected. I checked SXPMLBCWavefront.dat and found that it does have offsets of Z5 = -900 and Z6 = -110. These initial Z5 and Z6 offsets are discussed in IT 8513. But the initial Z6 offset is about 1 order of magnitude different from what is indicated in the comments & discussion. I thought that, last night, SX was removing a lot of astigmatism during the first round of FPIA, in the same sense as the offsets.

02:20 Both shutter doors are open! There is a mechanical problem, which Bruce, Mike & David resolved to help us open but it will need more work tomorrow.

2018 LB1 round 1

03:24 Slewed to copointing field for 2018 LB1 with TMS active. The field is low, 34 deg, but the others are also low or below 30 deg. Waiting for a TMS correction.

We stopped the TMS loop and ran dofpia, /x2. FPIA converged in 3 iterations, however the copointing images showed poor IQ on the blue side. We copointed, set the TMS reference.

Because of the poor IQ on LBCB, we repeated FPIA. On the blue side, z4 = 1200 was added. On the resulting pair of copointing images, the IQ was better, although the FWHM was about 9 pixels. This is consistent with the DIMM seeing.

04:00 We are starting to take data: NS, no guiding and (no TMS active; I forgot to start it until the last exposure as reading out). (I will fill in above the notes I lost). The PA is 0 on LBCB & 90 on LBCR.

2017 VW13 round 1

04:15 (about) Slewed to the copointing field. obs6 froze up - out of free memory. Opened the LBC UI on robs. I had TCSGUI, PMCGUIs, PSFGUIs up as well as NSGUI. I was displaying a png (display xxx.png), also had the FACSUM running as bos4.

04:20 Starting FPIA. It converged in 2 iterations but still with estimated seeing >2" on LBCB. Elevation xx. Set the TMS reference and started the loop. On the copointing images, 7.2 pix FWHM (1.6") on Blue and 6.7 pix FWHM on red.

04:31 Started the science observation. NS, TMS in active mode. This time it is guided as I forgot to uncheck the box. The guide star has FHWM ~2". On B 043228 there may be a tail.

04:41 Red guide star now has FWHM 1.2-1.4".

2019 XO2 round 1

04:46 Slewing with TMS active. Taking the pair of copointing images. Copointed. Checking IQ. FWHM 6.8 pix on Blue and 6 pix on Red. Stars are a bit out-of-round. We never had a very good collimation, so I'll try again. DIMM now reporting 1.5" (zen corr). There was some defocus needed. From the blue pupil, est seeing 1.52" and sending spherical corrections.

04:57 Repeating the CP images while setting the TMS reference, so we have some measure of IQ. On the pair of images, B 050000 and R 045946, the IQ is still not that good and FWHM is ~ 6.7 B and ~5.7 R

05:03 Starting the science. PA=0 on B and 90 on R. The guiding button was not checked, though I saw guiding failed on the UI and the resulting image looks like it has a jump. I think it was guiding.

05:18 Repeating the OB here, having checked and unchecked the guiding button. It was trying to guide again. I have dropped the exptime to 115 sec and will repeat (with guiding). The first pair of exposures is OK. Meanwhile, I made a copy of the script where I edited the value within the tag to be 0, not 1.

2018 LB1 round 2

05:31 Slewing with TMS active and wait for an update. Running dofpia to try to do better with collimation - images are out-of-round, but the main problem is the poor seeing. Elev 52 deg.

05:44 Copointing. w 05:48 Starting the science. This time it is not guiding. TMS active. IQ is now much better!, FWHM = 4.8 pix on red and blue copointing images.

05:55 obs6 came back to live. I found that the tms loop that I'd started from obs6 had not really exited by control-C. So I had two tms loops running between 04:15-05:55. I wonder how that was working, or not working.

2017 VW13 round 2

06:03 I took the pair of copointing images and assessed IQ. It was pretty good: 4.5pix Blue and 4.8 Red. Copointed

06:08 Started the science: NS, not guiding, TMS active. 60 deg elevation. DIMM 2" then 1.35" (zen-corr).

2019 XO2 round 2

06:19 Slewing to the copointing field - a long slew in azimuth.

06:22 The first TMS corrections came.

06:22 Taking the pair of copointing images. FWHM 4.5 on Blue and about 5.5 on Red (round). OK - we'll continue.

06:27 Starting the science: NS, not guiding and TMS active.

2018 LB1 round 3

06:42 Slewing to the copointing field and waiting for the first update.

06:44 Taking the pair of copointing images. The collimation on LBCR is poor.

06:47 Running FPIA. The first pupil images were converged according to it.

06:50 Taking a second pair of copointing images. FWHM 9 pix on Blue and 5.3 pix on Red. Skipping the copointing in the interest of time. We're not at very high elevation. This field is pointing to the S and the wind is from the E, hitting the left side.

06:52 Starting the science. NS, no guiding, TMS active.

2017 VW13 round 3

07:04 Slewing to copointing field.

07:06 Both sides got a TMS correction - taking the images. The IQ is ok: 6.2 pix on Blue and 5.5 on Red. Skipping copointing again.

07:09 Starting the science. NS, not guiding.

I noticed that sometimes the FACSUM displays the rates, but other times it doesn't.

2019 XO2 round 3

07:21 Slewing to the copointing field.

07:23 Slew completed - TMS update was made.

07:24 Taking the pair of copointing images. FWHM 6.7 pix on both Blue and Red. The images look round.

07:27 Starting the science. Trails on r 072731 look discontinuous, not smooth, but guiding is off and the LBTplot shows no guide corrections since 05:30 UT. The GUI reports that it is taking a guide exposure and then says "tracking disabled".

07:39 Finished

Summary

It was a rough start, and the seeing and collimation plagued the first round. Conditions were probalby not very stable and we had to recollimate a few times. The first target was just over 30 deg elevation and, in retrospect, it would have been better to collimate closer to the zenith and set the reference there. The second two rounds were better. Still the images had FWHM around 1.5" though sometimes down to about 1", but the IQ was holding and there was no need to recollimate. We noticed that Blue IQ was poor at times and Alex noted that for that object (one of the 12h ones) we were 90 deg from the wind, so there could have been dome seeing. Alex saw one or two moving objects when blinking a couple of images.

-- BarryRothberg - 04 Apr 2022
Topic revision: r16 - 08 Apr 2022, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback